find the slope of the line perpindicular to the line given by 3x-5y=8. show all work to get the answer
Answer by rothauserc(4718)

3x -5y = 8
:
first step is to solve this equation for y, this is standard slope intercept form, y = mx + b where m is the slope and b is the y intercept
:
5y + 8 = 3x
y = (3x/5) - (8/5)
:
the slope of the line perpendicular to 3x - 5y = 8 is the negative reciprocal of the slope(m) of this line
:
the slope of 3x -5y = 8 is (3/5), therefore
:
(-5/3) is the slope of the line perpendicular to 3x -5y = 8
: